Can you help me solve this problem.
Our topic is about quadratic equations by completing the square.
The product of two numbers is 10 less than 16 times the smaller.If twice the smaller is 5 more than the larger number, find the two numbers.

Answer by scott8148(6628)

"The product of two numbers is 10 less than 16 times the smaller" ___ S * L = 16S - 10

"twice the smaller is 5 more than the larger" ___ 2S = L + 5 ___ 2S - 5 = L

substituting ___ S(2S - 5) = 16S - 10 ___ 2S^2 - 21S + 10 = 0

2S^2 - 21S = -10 ___ S^2 - (21/2)S = -5

S^2 - (21/2)S + (-21/4)^2 = -5 + (-21/4)^2

(S - (21/4)^2 = (-80/16) + (441/16) = 361/16 = (±19/4)^2

S = (19/4) + (21/4) = 10 ___ or ___ S = (-19/4) + (21/4) = 1/2

substituting ___ 2(10) - 5 = L ___ L = 15 ___ or ___ L = 2(1/2) - 5 = -4


using the negative root causes the "larger" to be the "smaller" , and vice versa